3. Give an example of a predicate P(x,y) such that ∃x∀yP(x,y) and ∀y∃xP (x, y) have different truth values.


1
Expert's answer
2021-11-02T12:41:26-0400

Let us give an example of a predicate "P(x,y)" such that "\u2203x\u2200yP(x,y)" and "\u2200y\u2203xP (x, y)" have different truth values.

Let "P(x,y)" means "x\ge y," and the domains for "x" and "y" are the set "\N" of all positive integers.

Since the set "\N" does not contain the greatest element, the value of "∃x∀y(x\ge y)" is false. On the other hand, for any "y\in\N" let "x=y". Then "y\ge y=x," and hence the value of "\u2200y\u2203xP (x, y)" is true.